OpenStudy (anonymous):

Integrating factors... What do I use for my limits of my IF? I'll post where I am currently at. Problem is to solve for C in terms of t from 0 to t. Equation soon to come...

OpenStudy (anonymous):

\[\int\limits_{0}^{t}d[Ce^{\int\limits_{}^{}{\frac{ 6 }{ 50+6t }dt}}]=\int\limits_{0}^{t}{\frac{ 6 }{ 50+6t }}e^{\int\limits_{}^{}{\frac{ 6 }{ 50+6t }dt}}dt\]

OpenStudy (anonymous):

\[IF=e^{\int\limits\limits_{}^{}{\frac{ 6 }{ 50+6t }dt}}=e^{\ln{50+6t}}\]Right?

OpenStudy (anonymous):

\[e^{\ln{50+6t}}=50+6t\]\[\int\limits\limits_{0}^{t}d[C(50+6t)]=\int\limits\limits_{0}^{t}{\frac{ 6 }{ 50+6t }}(50+6t)dt\]

OpenStudy (anonymous):

@Zarkon @ganeshie8 @Luis_Rivera Anyone DE savvy and don't mind checking my setup steps?

OpenStudy (anonymous):

\[C(t)=\frac{ 15t }{ 50+6t }\]

OpenStudy (anonymous):

???

OpenStudy (anonymous):

there is a typo... line (1) needs to be a 15 in the numerator of the right hand side